A man is four times as old as his son. Three years ago, he was 5 times s old as his son was at that time. HOW OLD IS THE SON?

Answer by stanbon(75887)

A man is four times as old as his son. Three years ago, he was 5 times s old as his son was at that time. HOW OLD IS THE SON?
-----
Equations:
f = 2s
f-3 = 5(s-3)
-----
Substitute for "f" and solve for "s":
2s-3 = 5s-15
3s = 12
s = 4 yrs (the son's age now)
=============
Cheers,
Stan H.
==============